OpenStudy (alemace1):

I need help minimizing Q=3x^2+y^2, where x+y=4

OpenStudy (freckles):

write Q in terms of one variable using the equation x+y=4

OpenStudy (freckles):

is this a calculus or algebra question?

OpenStudy (alemace1):

Calc

OpenStudy (alemace1):

I'm confused

OpenStudy (freckles):

so you know to minimize or to maximize this involves finding the derivative of thing you wish to minimize or maximize ?

OpenStudy (alemace1):

Please help me step by step

OpenStudy (freckles):

Do you know how to use x+y=4 to write Q in terms of one variable ?

OpenStudy (freckles):

If not, here is a hint: Solve x+y=4 for one of the variables

OpenStudy (alemace1):

No i don't

OpenStudy (freckles):

There is no wrong or right variable to solve for in x+y=4 just pick one and solve for it

OpenStudy (alemace1):

Plug in any number ?

OpenStudy (freckles):

no solve x+y=4 for x or solve x+y=4 for y

OpenStudy (freckles):

don't do both just do one of those things

OpenStudy (freckles):

you only need one step to solve either equation

OpenStudy (freckles):

try subtracting something on both sides

OpenStudy (alemace1):

I'm lost

OpenStudy (freckles):

x+y=4 Say we want to solve for y notice you have a plus x next to it to undo an addition of x on that side with the y you will need to subtract x on both sides

OpenStudy (freckles):

what equation do we have after doing that ?

OpenStudy (freckles):

\[x+y=4 \\ (x+y)=4 \\ \\ (x+y)\color{red}{-x}=4\color{red}{-x} ... \\ \text{ what do we have after subtracting } \\ x \text{ on both sides } \\ \text{ what does the left hand side simplify to }\]

OpenStudy (alemace1):

Y=4-X

OpenStudy (freckles):

right now write Q in terms of x only

OpenStudy (alemace1):

Q=4((Y-6)Y+12) ??

OpenStudy (alemace1):

IS that the answer?

OpenStudy (freckles):

not entirely sure I see how you got that

OpenStudy (freckles):

didn't you say \[Q=3x^2+y^2 \text{ and } y=4-x\]

OpenStudy (alemace1):

Yeah

OpenStudy (freckles):

just replace y with (4-x) to write Q in terms of one variable

OpenStudy (alemace1):

So now I would be solving Q=3x^2+(4-x)^2 ?

OpenStudy (freckles):

You differentiate that and set Q'=0 and solve for x to find critical numbers

OpenStudy (alemace1):

So the x values are 0?

OpenStudy (freckles):

no solving Q'=0 for x gives the values of x for which the graph of Q has horizontal tangents and could possibly tell us where we have max or min if any

OpenStudy (alemace1):

How would I plug this into a graphing calculator?

OpenStudy (freckles):

Q is actually a parabola which is easy to find max/min of even without calculus this is why I asked if it was an algebra or calculus question but either way is easy i honestly like derivatives more :p parabolas have min/max occur at the vertex... and the way to determine if the vertex is a max/min is to determine if the parabola is open up or down and you can do this by looking at the coefficient of x^2 If this is a calculus question though, I suggest using calculus ways and differentiate Q you need power rule, constant rule, constant multiple rule, and you could also use chain rule if you don't expand that (4-x)^2 part

OpenStudy (freckles):

\[\frac{d}{dx}(3x^2)=? \\ \frac{d}{dx}(4-x)^2= ?\]

OpenStudy (alemace1):

Idk how to do this This is not helping